OpenStudy (anonymous):

Can someone please help me how to set this up? A police car recieves a radio call to catch a vehicle which is speeding down the highway at 80 mph. The police car, which is 12 miles away, drives after it at 108 mph. How long will it take for the police car to catch up?

OpenStudy (anonymous):

The car has a 12 mile head start, going 80 mph, so his distance is: \[d_{car} = 80*t+12\] The is going at 108 mph, so his distance is: \[d_{cop} = 108t\] Setting them equal to each other we get: \[80t+12 = 108t \Rightarrow 12 = 28t \Rightarrow t = \frac{12}{28} = \frac{3}{7}\] So 3/7 of an hour.

OpenStudy (anonymous):

about 25.7 mins.

OpenStudy (anonymous):

if there is something that needs more explaining just ask.